True or False.
1)To add two functions you simply add the corresponding y-coordinates to get the combined value.
2)When two functions are divided, the domain of the combined functions consists of all of the values in the domains of the original function.
3)If f(x) is a function that is defined for all xER, then f(f^inverse(x)) = x
4)To solve the inequality f(x) > g(x), a student could graph the combined function y=f(x)-g(x) and identify the portions of the graph that are below the x-axis.
5) When two functions are added,the domain of the combined functions consists of all of the values common to the domain of both of the original functions.
6) Determine the domain and range of the function y = sqrt(x^2+10x+24)

Answer by stanbon(75887)

True or False.
1)To add two functions you simply add the corresponding y-coordinates to get the combined value.
True
------------------
2)When two functions are divided, the domain of the combined functions consists of all of the values in the domains of the original function.
False
------------------
3)If f(x) is a function that is defined for all xER, then f(f^inverse(x)) = x
True
-----------------
4)To solve the inequality f(x) > g(x), a student could graph the combined function y=f(x)-g(x) and identify the portions of the graph that are below the x-axis.
False
-------------------
5) When two functions are added,the domain of the combined functions consists of all of the values common to the domain of both of the original functions.
True
-------------------------
6) Determine the domain and range of the function y = sqrt(x^2+10x+24)
To find the domain solve x^2+10x+24 >= 0
(x+6)(x+4) >= 0
True for (-inf,-6]U[-4,+inf)
That is the domain.
===========================
Cheers,
Stan H.